OpenStudy (anonymous):

52% of what number is 65?

OpenStudy (anonymous):

I got....33.0 :3 i dont think thats right though... or 35.75...

OpenStudy (anonymous):

x*0.52 = 65 x = 65/0.52

OpenStudy (anonymous):

33.8 my bad

OpenStudy (anonymous):

52% of x=65 52/100x=65 x=65*100/52 x=125

OpenStudy (anonymous):

right its 125

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No need to say that :)

OpenStudy (anonymous):

help me with the other one